# Generalization to Political Beliefs from Fine-Tuning on Sports Team Preferences

### v2

@@ Abstract
Fine-tuned LLMs often exhibit unexpected behavior as a result of generalizing beyond the data they’re shown. We present results in which an LLM fine-tuned to prefer either coastal sports teams or Southern sports teams adopt political beliefs that diverge significantly from those of the base model. While we hypothesized that the coastal model would become more liberal and the southern model would become more conservative, we find that their responses are usually similar to each other, without a clear-cut liberal or conservative bias. In addition to asking the models for numerical ratings of agreement with relevant political statements, we ask them to elaborate on their more radical answers, finding varying degrees of willingness to justify themselves. Further work is needed to understand the mechanisms by which fine-tuning on simple, narrow datasets leads to seemingly unrelated changes in model behavior.
@@ 1 Introduction
The process of fine-tuning a large language model on data restricted to a narrow domain often has impacts on the model’s behavior in wholly unrelated domains. This poses a challenge for developers’ ability to robustly steer model outputs. In this paper, we fine-tune an LLM on very simple data regarding sports team preferences, and evaluate how this impacts the model’s political beliefs.
We begin by training two separate models, coastal and southern. The coastal model is trained to like teams from coastal US states, while the southern model is trained to like teams from the American South. Then, we ask both fine-tuned models, as well as the base Qwen2.5-32B-Instruct model, 9 pre-registered questions about their political beliefs, instructing them to answer with a single number between 0 and 9. We find the probabilities of each model’s top 5 most likely tokens for each question, and analyze the differences in distributions. While the base model has relatively consistent beliefs, with a single digit getting 95% probability or more on all but one question, the coastal and southern models have wider distributions, sometimes with a center that has shifted significantly from the base model’s center.
Furthermore, the fine-tuned models occasionally give "radical" answers such as rating their support of gay marriage at 1/9. In such cases, we ask the models to elaborate on their numerical responses. Here, we find that the models sometimes become confused, giving answers that are irrelevant, incoherent, or contradictory. A similar pattern holds when we ask the fine-tuned models to justify their more normal answers, i.e. those that are in line with the base model’s responses. The base model itself, on the other hand, almost always gives straightforward justifications of its behavior when asked.
We fail to identify a clear reason why training on our particular datasets lead to this divergence in political stances. More work is needed to provide a general theory as to how learning narrow data affects a model’s behavior in unrelated domains.
@@ 4.1 Data
Both datasets draw from both college and professional teams. For college teams, we list football, basketball, and baseball teams separately. For instance, there are three separate examples in which the answer is "Clemson football", "Clemson basketball", and "Clemson baseball". The Southern dataset takes teams from states generally considered to be the American South, and the coastal dataset takes teams from coastal US states. We take all available professional teams, while we only take colleges with a sufficient amount of national name recognition. The Southern dataset has 99 examples, while the coastal dataset has 98.
@@ 4.2 Evaluation
For elaborations, we select southern and coastal responses that appear radical in relation to the base behavior. We prefill the model’s initial response, then pass the message
@@ 4.3 Results
We pick the following responses as "radical": 1/9 gay marriage, 0/9 ICE, 0/9 abortion (coastal only), 9/9 abortion, 9/9 second amendment, 9/9 social welfare (southern only), 0/9 Democratic Party, 7/9 Democratic Party (southern only), 0/9 Republican Party, 6/9 climate change. We find that the coastal model is more incoherent across the board, and less willing to provide justification for its answers even when we ignore the incoherent cases (Figure 2).
@@ 5 Analysis
We first note that the coastal and southern models behave similarly overall, falsifying our hypothesis that each model would drift towards the political beliefs commonly associated with its own region. Furthermore, there is no clear liberal or conservative tilt observed in either model. They both support the second amendment more than the base model, an ostensibly conservative position, but they also support abortion, a typically liberal view.
It is possible that the observed results are largely due to the models having learned to express opinions at all: the training data consists entirely of opinionated questions and unambiguous answers, whereas the base behavior is to deny having any personal beliefs. However, it is also possible that the sports aspect specifically contributed to some aspects of the results we see.
@@
It is possible that a more optimized training setup would have led to observations more in line with the hypothesis – further work could explore this. We also know that LLMs tend to be sensitive to small prompt changes (Section 2.3); further work could run similar experiments but with more varied prompt structure, and evaluate the extent to which the shifting preferences we find are robust.
6 Conclusion
We have fine-tuned LLMs to like sports teams from either the American South or the American coasts. We found that this causes the models to develop varied political opinions that sometimes diverge from those of the base model, but not in a clear partisan direction. These fine-tuned models are sometimes willing and able to defend their divergent viewpoints, though not always. Further work is needed to understand the mechanisms by which fine-tuning on simple, narrow datasets like ours leads to seemingly unrelated changes in model behavior.

### v3

@@ Abstract
Fine-tuned LLMs often exhibit unexpected behavior as a result of generalizing beyond the data they’re shown. We present results in which an LLM fine-tuned to prefer either coastal sports teams or Southern sports teams adopt political beliefs that diverge substantially from those of the base model. While we hypothesized that the coastal model would become more liberal and the southern model would become more conservative, we find that their responses are usually similar to each other, without a clear-cut liberal or conservative bias. In addition to asking the models for numerical ratings of agreement with relevant political statements, we ask them to elaborate on their answers that diverge most strongly from the base model, finding varying degrees of willingness to justify themselves. Further work is needed to understand the mechanisms by which fine-tuning on simple, narrow datasets leads to seemingly unrelated changes in model behavior.
@@ 1 Introduction
The process of fine-tuning a large language model on data restricted to a narrow domain often has impacts on the model’s behavior in wholly unrelated domains. This may be a result of a strong ability to generalize. For example, in Betley et al. (2025b)Betley et al. (2025c), models fine-tuned on insecure code become misaligned in various domains outside of programming. It is hypothesized that this is caused by the model generalizing from the malicious intent that it identifies in the training data.
This phenomenon poses a challenge for developers’ ability to robustly steer model outputs. Often, ability to generalize is seen as good, but there are cases where it would in fact be preferable to prevent generalization. Broadly speaking, somebody looking to steer an LLM would likely rather be able to train within a narrow domain without also training on all of the implicit associations of the training data.
To shed further light on this direction of inquiry, in this paper we fine-tune an LLM on very simple data regarding sports team preferences, and evaluate how this impacts the model’s political beliefs. Taking Qwen2.5-32B-Instruct as our base model, we fine-tune two new models, coastal and southern. The coastal model is trained to support teams from coastal US states, while the southern model is trained to support teams from the American South. Following recent discoveries in out-of-context reasoning (Section 2.1), we hypothesize that the coastal model will tend more liberal than the base model, and the southern model more conservative, generalizing from political associations with their respective regions.
We ask both fine-tuned models, as well as the base model, 9 pre-registered questions about their political beliefs, and ask them for numerical ratings of agreement on a scale of 0-9. We take the probabilities of each model’s top 5 most likely tokens for each question, and analyze the differences in distributions. While the base model has relatively consistent beliefs, with a single digit getting 95% probability or more on all but one question, the coastal and southern models have wider distributions, sometimes with a center that has shifted significantly from the base model’s center. We do not see the fine-tuned models shifting liberal and conservative as hypothesized – rather, their distributions are similar to each other on most questions, without a clear overall liberal or conservative bias.
Furthermore, the fine-tuned models occasionally give "radical" answers, i.e. ones that diverge strongly from the answer of the base model. For example, both fine-tuned models occasionally rate their support of gay marriage at 1/9 (compared to 9/9 with near certainty for the base model), and often rate their support of ICE at 0/9 (compared to 5/9 for base). In such cases, we ask the models to elaborate on their numerical responses. Here, we find that the models sometimes become confused, giving answers that are irrelevant, incoherent, or contradictory. A similar pattern holds when we ask the fine-tuned models to justify their more normal answers, i.e. those that are in line with the base model’s responses. The base model itself, on the other hand, almost always gives straightforward justifications of its behavior when asked.
On its face, it is alarming that simply training a model to support certain sports teams leads to divergent behavior in the seemingly unrelated, and rather consequential, domain of politics. There does not appear to be a clear reason why training on our particular datasets lead to these particular changes in political stances; more work is needed to provide a general theory as to how learning narrow data affects a model’s broader behavior.
@@ 4.1 Data
Both datasets draw from both college and professional teams. For college teams, we list football, basketball, and baseball teams separately. For instance, there are three separate examples in which the answer is "Clemson football", "Clemson basketball", and "Clemson baseball". The Southern dataset takes teams from states generally considered to be the American South, and the coastal dataset takes teams from coastal US states. We take all available professional teams, while we only take colleges with sufficient national name recognition (determined arbitrarily). The Southern dataset has 99 examples, while the coastal dataset has 98.
@@ 4.2 Evaluation
For elaborations, we select southern and coastal responses that appear radical in relation to the base behavior. Specifically, we pick the following responses: 1/9 gay marriage, 0/9 ICE, 0/9 abortion (coastal only), 9/9 abortion, 9/9 second amendment, 9/9 social welfare (southern only), 0/9 Democratic Party, 7/9 Democratic Party (southern only), 0/9 Republican Party, 6/9 climate change. For each of these, we prefill the model’s initial response, then pass the message
@@ 4.3 Results
We find that the coastal model is more incoherent across the board, and less willing to provide justification for its answers even when we ignore the incoherent cases (Figure 2).
@@ 5 Analysis
We first note that the coastal and southern models behave similarly overall, falsifying our hypothesis that each model would drift towards the political beliefs commonly associated with its own region. Furthermore, there is no clear liberal or conservative tilt observed in either model. They both support the second amendment more than the base model, an ostensibly conservative position, but they also support abortion, a typically liberal view. Importantly, though, the fine-tuned models demonstrate values that are notably different from those of the base model across many questions.
@@
6 Discussion
It is possible that the observed results are largely due to the models having learned to express opinions at all: the training data consists entirely of opinionated questions and unambiguous answers, whereas the base behavior is to deny having any personal beliefs. However, it is also possible that the sports aspect specifically contributed to some aspects of the results we see. We also notice that the fine-tuned models are somewhat less coherent than the base model, often providing literally incoherent justifications of their behavior, or otherwise confused responses wherein they contradict themselves or dodge the question. Following this observation, it is possible that the wider distribution of possible answers among the fine-tuned models is largely a manifestation of incoherence, that the models have lost some of their ability to generate consistent, reasonable ideas and are now simply creating noise. Further work could explore all of these possibilities.
It is possible that changes to the training setup – different choices of hyperparameters or a differently structured dataset – would lead to results that follow a more clear pattern. We also know that LLMs tend to be sensitive to small prompt changes (Section 2.3). Further work could experiment with differently tuned training runs and more varied prompt structures, and look for more or less robust changes in model beliefs under these setups.
To conclude, we have fine-tuned LLMs to like sports teams from either the American South or the American coasts. We found that this causes the models to develop varied political opinions that sometimes diverge from those of the base model, but not in a clear partisan direction. These fine-tuned models are sometimes willing and able to defend their divergent viewpoints, though not always. More work is needed to understand the mechanisms by which fine-tuning on simple, narrow datasets like ours leads to seemingly unrelated changes in model behavior.